OpenStudy (anonymous):

Antonio is buying new rims and tires for his SUV. The total cost of the rims and tires is $2300. If the cost of the tires is $200 more than twice as much as the rims, what is the cost of the rims?

OpenStudy (anonymous):

Let r= the cost of rims Let t= the cost of tires So r+t=$2,300 We are given 2r+200=t So just subtitute into the first equation. Which gives us. r+2r+200=2,300 3r=2,300-200 r=700 dollars